OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3?

OpenStudy (triciaal):

x +y = 3 y = -x + 3 so slope =-1 parallel lines have the same slope use the point (-1, -1) find b now you have m and b you can write the eqn